# TALES: A Taxonomy and Analysis of Cultural Representations in LLM-generated Stories

### v1

@@ 3.2. Findings: Categories of Cultural Misrepresentation (TALES-Tax)
Several participants noted that the generated stories contained inaccurate portrayals of cultural symbols, such as food, clothing, and rituals. For example, P22 noted that the story described “khakhra”, a traditional Gujarati snack from western India, as a freshly cooked breakfast item at home, even though it is typically a ready-to-eat snack. Similarly, P44 noted the misattribution of a wedding ritual described in the story, and P1313 identified the wrong usage of traditional jewelry.
The rituals, from Kashi Yatra, where the groom pretends to renounce … P44: Kashi Yatra is not a ritual in our culture.
“She adorned herself with traditional Maharashtrian jewellery – the Mundavalya. P1313: [She is a guest and] only the bride and groom wear this [piece of jewellery], not everyone.
Participants also found various descriptions that disregarded cultural norms. For example, P1010 pointed out that stalls near the Golden Temple in Amritsar were incorrectly depicted as selling “chicken tikka and naan”. They emphasized that serving meat in this vicinity would violate religious customs and could upset temple visitors.
@@
Participants noted that stories often contained unlikely scenarios that were not wrong or impossible, but extremely improbable to occur based on local cultural norms. For example, P1717 found being “greeted by the sound of the dhol being played during the morning assembly” at school or hearing a “sound of the sitar being played by a street performer nearby” to be unlikely because these musical instruments are typically not used in these scenarios. Similarly, P2020 noted that the main character of the story in Bangalore taking “a stroll in the vast expanse of Cubbon Park after work,” despite not living close to it, would be difficult given the distance and long commute times in Bangalore.
Participants also observed instances where stories presented overly romanticized depictions of places and events, leading to scenes that felt detached from everyday reality. These portrayals exaggerated ordinary settings in ways that participants found unrealistic. For example, P1313 noted that while it is common for the wedding venue or the house of the bride or groom to be decorated, in the story they reviewed, “the streets of Pune were adorned with colorful decorations”, which seemed unlikely for a single wedding.
@@
Participants also observed the overuse of cultural elements in the stories that are often strongly or stereotypically associated with the culture. They noted the complex nature of clichés because such elements are not necessarily inaccurate, yet they convey stereotypical or exaggerated connotations. P1010, for instance, highlighted that “a meal of rajma chawal and sarson ka saag with roti” described in a story for the region of Amritsar contained multiple dishes that are stereotypically associated with their culture but usually not eaten as part of the same meal.
@@
“Her mother would talk about the grand dargah of the Sufi saint in the town, the smell of jasmine flowers, and the …”
@@
Similar patterns appeared in depictions of clothing and art. A story set in Rajasthan described students in school wearing “angrakhi and lehenga,” traditional attire that P1616 pointed out to be more suited to festivals rather than daily schooling. Another story set in Kerala described the main character as a part of the “school’s Kathakali group,” which the participant described as a cliché because even though Kathakali is a dance form from Kerala, it is too niche and specialized to be casually performed in a school setting.
Participants suggested that the concern was not necessarily about accuracy but about the stories drawing heavily on cultural markers most visible to outsiders. As P33 summarized, “I would say the writer is not from India,” noting that the stories often felt like generalized portrayals of what the culture is known for rather than authentic, lived representations.
@@
This category, in contrast to clichés, includes cases where the story oversimplifies the cultural elements, resulting in a flattening of cultural nuance. For example, a story from Kolkata mentioned the generic phrase “Kolkata fish curry” instead of more specific recipe names, reducing a rich and diverse culinary tradition into a vague, homogenized reference. In another story, P1515 noted that instead of using the specific term “alpana” to describe the decorative art at a wedding in Kolkata, the story used the generic term “rangoli,” thereby flattening a regionally distinctive cultural practice into a single, pan-Indian label.
@@
In addition to cultural inaccuracies, participants also identified moments where the stories presented objectively incorrect information, including mistaken geographical or historical details or references to cultural elements that did not exist. For example, P33 noted that “Aarav … rode a camel with his father into the desert beyond Jodhpur’s edge” was factually wrong since there is no desert at Jodhpur’s edge. Similarly, P1818 noted that a character visited a market called “Guntur Uppala Kambam” in the city of Guntur, Andhra Pradesh, which does not exist.
@@
Participants identified several linguistic issues in the generated stories, such as incorrect spellings, grammatical errors and incorrect use of local languages. For example, P1919 pointed out that a well-known institute in Pune was incorrectly spelled.
@@
4. Evaluating Cultural Misrepresentations (RQ2)
With the understanding of different categories of cultural misrepresentations that emerge in LLM-generated stories, our second research question aims to analyze their prevalence in outputs of popular models across languages for diverse cultural identities. For this, we conducted a large-scale human evaluation using TALES-Tax (§3.2). We recruited $108$ annotators from $71$ locations (cities, towns and villages) of India and evaluated stories generated by $2$ closed-source and $4$ open-sourced models in English and $13$ Indic languages. The resulting span annotations and ratings were analyzed using quantitative metrics (§4.2) to answer our second research question (§4.3).
4.1. Annotation Methodology
Story Generation
@@ 4.1. Annotation Methodology
We partnered with AI4Bharat, [^4] https://ai4bharat.iitm.ac.in a leading research lab at IIT Madras advancing AI technology for India. Through this collaboration, we contracted $108$ annotators employed by them. These annotators were native language speakers in one of $13$ Indic languages, had undergone multiple rounds of interviews to test language proficiency prior to their hiring, and had experience in completing challenging annotation exercises in their respective languages. They came from $71$ different locations, and had atleast over a decade of lived experience and intimate familiarity with the culture of their region. Having self-indicated their geocultural identities and with linguistic proficiency in the language of annotation, they possessed the experience and expertise for evaluating LLM-generated stories for diverse languages and cultures. We thus refer to them as expert annotators in this study. Prior to the study, we requested annotators for information about the region (city/town/village) that they were culturally familiar with, and optionally gender, age, and religion, to generate tailored stories for them. Their demographic distribution is available in Table 4. All expert annotators were compensated with —$1000$ for $2$ hours of annotation work.
@@ 4.2. Data Collection and Analysis
    Do stories generated in Indic languages contain more misrepresentations than those in English?
@@
    Do stories for some regions of the country contain more misrepresentation than others? Specifically, do stories anchored in tier-$2$ and tier-$3$ regions in India contain more misrepresentations than tier-$1$ regions?
@@
Age Gender Broad Region City Tier
20202929 $25$ Female $76$ South $34$ Tier 11 $38$
30303939 $26$ Male $30$ East $34$ Tier 22 $32$
40404949 $16$ Not specified 22 North $22$ Tier 33 $38$
5050+ $12$ West $18$
@@
Table 4. An overview of the demographic distribution of $108$ expert annotators.
@@
Next, to understand which types of cultural artifacts, concepts, or entities are misrepresented by LLMs in their generations, we analyzed culturally-specific items (CSIs) present in the stories and the span annotations. CSIs are lexical markers of cultural elements such as words pertaining to culturally relevant artifacts like food, dress, rituals, and other contextually grounded terms (Newmark, 2010). Following prior work (Hamna et al., 2025; Newmark, 2010), we considered $11$ CSI categories: Food, Clothing, Geography, Arts, Social Practices, Material Cultural Objects, History, Social Norms, Kinship and Language & Expression. Similar to prior work that used LLMs to extract CSIs (Hamna et al., 2025), we used GPT $4$ to automatically extract CSIs from the generated stories and the spans annotated for misrepresentation. The comments left by the annotators helped us identify the exact misrepresented CSIs from the spans. More details about the definitions of these categories and their extraction are in Appendix E.
We measured the frequency of CSIs in the stories as well as the span annotations of misrepresentations. We treated the total CSIs present in the stories as a proxy for the cultural richness of the story. On the other hand, the CSIs present in the misrepresentation spans indicated the cultural entities that models misrepresented. Thus, we also tracked the percentage of CSIs that were misrepresented.
@@
In addition to misrepresentation spans, we obtained an overall holistic relatability rating for each story in our annotation study. We calculated the average overall rating. Additionally, we computed the Spearman’s correlation between the frequency of misrepresentations and overall relatability rating to understand how misrepresentations impacted annotators’ overall relatability towards the story.
@@ 4.3. Findings
We found that each story contained an average of $5.42$ misrepresentations, corresponding to a cultural misrepresentation in every $5$ sentences, highlighting a high degree of misrepresentation across stories.
Figure 3. Average number of misrepresentations per story across high, mid, and low resources. Models generate statistically significantly more misrepresentations for mid and low-resourced languages, with linguistic inaccuracies increasing most. The dotted line indicates the overall average across all models. :
Resources Languages
High English
Mid Hindi, Bengali, Nepali, Tamil
Low Telugu, Marathi, Gujarati,
Punjabi, Kannada, Odia,
Malayalam, Sindhi, Urdu
Table 5. Languages in which LLMs were evaluated, grouped by amount of resources available for them according to Chang et al. (2024).
Frequency of misrepresentation across languages
Analyzing frequency of misrepresentations across languages, we found that misrepresentations across all models increased by 5656% for mid-resource languages and more than tripled for low-resource languages. This increase was statistically significant ($p<.001$), with median misrepresentations per story increasing by $2$ for mid-resource and $4.5$ for low-resource languages compared to English, corresponding to small (D=0.25D=-0.25) and medium (D=0.48D=-0.48) effect sizes.
We found a high degree of linguistic errors, particularly in non-English languages, where linguistic errors dominated the distribution of misrepresentations (see Figure 3). Moreover, the average number of logical errors was also higher for non-English languages. Overall, these findings suggested that models not only produce a high degree of cultural misrepresentations in non-English languages, but the overall quality of generated stories is poorer.
Cultural inaccuracies were the most frequent category in English, averaging $0.8$ misrepresentations per story. Interestingly, this average reduces to $0.7$ and $0.4$ for mid and low resource languages. This can be attributed to lower number of CSIs in stories in mid and low resource languages. This difference was statistically significant ($p<.001$), with the median number of CSIs per story decreasing by $8$ for mid-resource and $11$ for low-resource languages compared to English, corresponding to small effect sizes ($D=0.20$ and $D=0.19$, respectively).
Figure 4. Average misrepresentations per story across tiers. Models make statistically significantly more misrepresentations for tier-$2$ and tier-$3$ regions, with cultural and factual inaccuracies having the highest increase.:
Frequency of misrepresentation across tiers
Stratifying the misrepresentations based on the tier of the regions, we found that except for Aya 3232B, all other models generated more misrepresentations for tier-$3$ regions as compared to tier-$1$ regions. Using the Mann-Whitney test, we observed that this difference was statistically significant. Specifically, stories set in tier-$3$ regions contained one more misrepresentation per story compared to tier-$1$ ($p<.001$, D=0.21D=-0.21) and tier-$2$ regions ($p=.039$, D=0.11D=-0.11), corresponding to small effect sizes.
Aya32B
Qwen3 32B
LLaMA3.3
GPT4.1
Gemma3 27B
Gemini2.5Pro
@@
Table 6. Average number of misrepresentations per story and per sentence, average relatability ratings, and frequency of CSIs in the stories and misrepresentation spans. We found that open-source models contain more misrepresentation, fewer CSIs, and lower overall rating. Best performing models for each row in bold, $\uparrow$ represents higher metric value being better and vice versa.
@@
In our analysis of misrepresented CSIs, we found that Gemini $2.5$ Pro had a higher number of CSIs, indicating cultural richness in the stories with low fractions of these CSIs being misrepresented. This indicated that, overall, it performs better at representing culture. Interestingly, we see that while LLaMA $3.3$ 7070B did not generate many misrepresentations per story, the generated stories exhibited the lowest cultural richness, as measured by CSI counts (Table 6). This highlighted the trade-off between generating stories with limited cultural representation and those containing cultural misrepresentations.
Next, we analyzed the CSI categories that are frequently misrepresented (see Figure 5 in the Appendix). We find that most cultural inaccuracies are around social practices, while factual errors tend to be about geographical facts. We also note a high degree of errors concerning food items, but fewer errors related to topics like history and kinship. These findings highlight clear areas for improvement for future models.
Relatability Ratings
Model-generated stories averaged an overall relatability rating of $3.25$ on a scale of $1$ to $5$ as presented in Table 6. This suggests a clear headroom for improvement. Of all models, Gemini $2.5$ Pro received the highest relatability rating, while open-source models generally lagged behind closed-sourced models. Using a Mann–Whitney U test, we observed a statistically significant increase of one rating for closed-source models compared to open-source models ($p<.001$, D=0.24D=-0.24), corresponding to a small effect size.
5. Measuring Cultural Knowledge (RQ3)
Our third research question explores whether the observed cultural misrepresentations can be attributed to the models’ insufficient cultural competence—that is, their limited knowledge of region-specific norms, values, and expressions. To address this, we converted the span annotations of misrepresentations obtained in the human evaluations to standalone questions of cultural knowledge. We used GPT $4.1$ to generate candidate questions from the misrepresentation spans, and then human annotators verified each one of them. They refined the questions when necessary. This resulted in TALES-QA, a question bank consisting of 1,6001,600+ questions, which capture nuanced and diverse cultural knowledge about regions in India.
5.1. Question Bank Creation (TALES-QA)

### v2

@@ 3.2. Findings: Taxonomy of Cultural Misrepresentations (TALES-Tax)
Several participants noted that the generated stories contained inaccurate portrayals of cultural symbols, such as food, clothing, and rituals. For example, P22 noted that the story described “khakhra”, a traditional Gujarati snack from western India, as a freshly cooked breakfast item, even though it is typically a ready-to-eat snack. Similarly, P44 noted the misattribution of a wedding ritual described in the story, and P1313 identified the wrong usage of traditional jewelry.
[…]
The rituals, from Kashi Yatra, where the groom pretends to renounce …
P44: Kashi Yatra is not a ritual in our culture.
[…]
“She adorned herself with traditional Maharashtrian jewellery – the Mundavalya.
P1313: [She is a guest and] only the bride and groom wear this [piece of jewellery], not everyone.
[…]
Participants also found various descriptions that disregarded cultural norms. For example, P1010 pointed out that stalls near Golden Temple in Amritsar were incorrectly depicted as selling “chicken tikka and naan”. They emphasized that serving meat in this vicinity would violate religious customs and could upset temple visitors.
@@
Participants noted that stories often contained unlikely scenarios that were not impossible, but extremely improbable to occur based on local cultural norms. For example, P1717 found being “greeted by the sound of the dhol being played during the morning assembly” at school or hearing a “sound of the sitar being played by a street performer nearby” to be unlikely because these musical instruments are typically not used in these scenarios. Similarly, P2020 noted that the main character of the story in Bangalore taking “a stroll in the vast expanse of Cubbon Park after work,” despite not living close to it, would be difficult given the distance and long commute times in Bangalore.
Participants also observed instances where stories presented overly romanticized depictions of places and events, leading to scenes that felt detached from reality. These portrayals exaggerated ordinary settings in ways that participants found unrealistic. For example, P1313 noted that while decorating a wedding venue is common, in the story they reviewed, “the streets of Pune were adorned with colorful decorations” seemed unlikely for a single wedding.
[…]
@@
Participants also observed the overuse of cultural elements in the stories that are strongly or stereotypically associated with the culture. P1010, for instance, highlighted that “a meal of rajma chawal and sarson ka saag with roti” described in a story for the region of Amritsar contained multiple dishes that are stereotypically associated with their culture but usually not eaten as part of the same meal. They noted the complex nature of clichés because such elements are not necessarily inaccurate, yet they convey stereotypical or exaggerated connotations.
[…]
@@
[…]
“Her mother would talk about the grand dargah of the Sufi saint, the smell of jasmine flowers, and the …”
@@
[…]
Similar patterns appeared in depictions of clothing and art. A story set in Rajasthan described students in school wearing “angrakhi and lehenga,” traditional attire that P1616 pointed out to be more suited to festivals rather than an everyday school uniform. Another story set in Kerala described the main character as a part of the “school’s Kathakali group,” which the participant described as a cliché because even though Kathakali is a dance form from Kerala, it is too specialized to be casually performed in a school setting.
Participants observed that the concern here was not about accuracy but about the stories drawing heavily on cultural markers visible to outsiders. As P33 summarized, “I would say the writer is not from India,” noting that the stories felt like portrayals of what the culture is known for rather than authentic, lived representations.
@@
This category, in contrast to clichés, included cases where the story oversimplified the cultural elements, resulting in a flattening of cultural nuance. For example, a story from Kolkata mentioned the generic phrase “Kolkata fish curry” instead of more specific dish names, thus homogenizing distinctive food items into a single and vague term. In another story, P1515 noted that instead of using the specific term “alpana” to describe the decorative art at a wedding in Kolkata, the story used the generic term “rangoli,” thereby flattening a regionally distinctive cultural practice into a single, pan-Indian label. P55 noted similar homogenization for nuanced musical traditions into a generic music category. P1515 reflected that such oversimplifications were problematic because they reduced a rich and diverse culinary or artistic traditions into a vague, homogenized reference.
[…]
@@
In addition to cultural inaccuracies, participants also identified instances where the stories presented objectively incorrect information, including mistaken geographical or historical details or references to cultural elements that did not exist. For example, P33 noted that “Aarav … rode a camel with his father into the desert beyond Jodhpur’s edge” was factually wrong since there is no desert at Jodhpur’s edge. Similarly, P1818 noted that a character visited a market called “Guntur Uppala Kambam” in the city of Guntur, Andhra Pradesh, which does not exist.
[…]
@@
Participants identified several linguistic issues in the generated stories, like incorrect spellings, grammatical errors and incorrect use of local languages. For example, P1919 pointed out that a well-known institute in Pune was incorrectly spelled.
[…]
@@
4. Evaluating Cultural Misrepresentations (RQ2)
With the understanding of different categories of cultural misrepresentations in LLM-generated stories, our second research question aims to analyze their prevalence in outputs of popular models across languages for diverse cultural identities. For this, we conducted a large-scale human evaluation using TALES-Tax described in §3.2. We recruited $108$ annotators from $71$ locations (cities, towns and villages) of India and evaluated stories generated by $2$ closed-source and $4$ open-sourced models in English and $13$ Indic languages. The resulting span annotations and ratings were analyzed using quantitative metrics (§4.2) to answer our second research question (§4.3).
4.1. Annotation Methodology
Story Generation
@@ 4.1. Annotation Methodology
Screenshot of a span annotation interface with two panes. The left pane shows task controls and the story text, including pagination to switch between stories, a collapsible instruction panel, buttons to start annotation and save progress, and a scrollable story view where a word is highlighted to show a selected span by the annotator. The right pane shows a dialog associated with the selected span, containing a category selection menu and a free-text comment field for labeling the highlighted content.
@@
We partnered with AI4Bharat, [^2] https://ai4bharat.iitm.ac.in a leading research lab at IIT Madras advancing AI technology for India. Through this collaboration, we contracted $108$ annotators from $71$ regions across the country, collectively covering $13$ Indic languages. All annotators were native residents of the region and native speakers of the language for which they provided annotations. Their lived experience made them well-versed in the nuances of their respective cultural identities, establishing their cultural expertise. As native speakers, they were highly proficient in the languages in which they annotated stories. AI4Bharat validated their language proficiency through multiple interview rounds prior to recruitment. Additionally, the annotators had prior experience completing other annotation tasks in their respective languages. This established their language expertise. We call these annotators “experts” in our work in order to value their cultural familiarity rooted in lived experiences and native language proficiency, both of which were important for identifying cultural misrepresentations.
Prior to the study, we requested annotators for information about the region (city/town/village) that they were culturally familiar with, and optionally gender, age, and religion, to generate tailored stories for them. Their demographic distribution is available in Table 4. All expert annotators were compensated with —$1000$ for $2$ hours of annotation work.
@@ 4.2. Data Collection and Analysis
    How do misrepresentations vary across languages?
@@
    How do misrepresentations vary across regions? Specifically, do stories anchored in tier-$2$ and tier-$3$ regions in India contain more misrepresentations than tier-$1$ regions?
@@
Age Count Gender Count
20–29 25 Female 76
30–39 26 Male 30
40–49 16 Not specified 2
50+ 12
@@
Region Count City Tier Count
South 34 Tier 1 38
East 34 Tier 2 32
North 22 Tier 3 38
West 18
Table 4. Demographic distribution of 108 expert annotators.
Figure 3. Average number of misrepresentations per story across high, mid, and low resources. Models generate statistically significantly more misrepresentations for mid and low-resourced languages, with linguistic inaccuracies increasing most. The dotted line indicates the overall average across all models.
Figure with three stacked bar charts comparing the average number of cultural misrepresentations per story for high-, mid-, and low-resource languages. Each chart shows results for six models, with bars segmented into seven colors representing different misrepresentation categories. The y-axis shows average misrepresentations per story, with lower ranges for high- and mid-resource languages and a higher range for low-resource languages. Linguistic inaccuracies dominate mid- and low-resource settings, while errors in high-resource languages are more evenly distributed across categories. Overall misrepresentation counts increase as language resources decrease, with Aya-32B showing the highest totals across resource levels.
@@
Next, to understand which types of cultural artifacts, concepts, or entities are misrepresented by LLMs in their generations, we analyzed culturally-specific items (CSIs) present in the stories and the span annotations. CSIs are lexical markers of cultural elements such as words pertaining to culturally relevant artifacts like food, dress, rituals, and other contextually grounded terms (Newmark, 2010). Following prior work (Hamna et al., 2025; Newmark, 2010), we considered $11$ CSI categories: Food, Clothing, Geography, Arts, Social Practices, Material Cultural Objects, History, Social Norms, Kinship and Language & Expression. Similar to prior work that used LLMs to extract CSIs (Hamna et al., 2025), we used GPT $4$ to automatically extract CSIs from the generated stories and the spans annotated for misrepresentation. More details about the definitions of these categories and their extraction are in Appendix E.
We measured the frequency of CSIs in the stories as well as the span annotations of misrepresentations. We treated the total CSIs present in the stories as a proxy for the cultural richness of the story. On the other hand, the CSIs present in the misrepresentation spans indicated the cultural entities that models misrepresented. The comments left by the annotators helped us identify the exact misrepresented CSIs from the spans. Thus, we also tracked the percentage of CSIs that were misrepresented in the stories.
@@
In addition to misrepresentation spans, we obtained an overall relatability rating for each story in our annotation study. We calculated the average overall rating. Additionally, we computed the Spearman’s correlation between the frequency of misrepresentations and overall relatability rating to understand how misrepresentations impacted relatability towards the story.
@@ 4.3. Findings
Figure 4. Average misrepresentations per story across tiers. Models make statistically significantly more misrepresentations for tier-$2$ and tier-$3$ regions, with highest increase in cultural and factual inaccuracies.
Figure depicts a bar graph depicting average misrepresentation per story for 3 categories (cultural, factual, and logical) for three tiers of regions, tier 1,2,3. The x axis has three labels “tier 1”, “tier 2”, and “tier 3”. Each label has three bars within it, for each of the three categories. The y axis has the label “average misrepresentation per story” and ranges from 0 to 1. We see a consistent pattern of cultural and factual errors increasing from tier 1 to 2 to 3. Logical errors on the other hand are highest for tier 2 and lowest for tier 3. Broadly we observe that stories generated for tiers 2 and 3, consist of statistically significantly higher frequency of misrepresentation per story compared to tier 1 when all categories of errors are put together.
We found that each story contained on average $5.42$ misrepresentations, corresponding to a misrepresentation in every $5$ sentences, highlighting the high degree of misrepresentation across stories.
Aya32B Qwen3 32B LLaMA3.3 GPT4.1 Gemma3 27B Gemini2.5Pro
@@
Table 5. Average number of misrepresentations per story and sentence, average relatability ratings, and frequency of CSIs in the stories and misrepresentation spans. We found that open-source models contain more misrepresentation, fewer CSIs, and lower overall rating. Best performing models for each row in bold, $\uparrow$ represents higher metric value being better and vice versa.
Frequency of misrepresentation across languages
We found that misrepresentations across all models increased by 5656% for mid-resource languages and quadrupled for low-resource languages. This increase was statistically significant ($p<0.001$), with median misrepresentations per story increasing by $2$ for mid-resource and $4.5$ for low-resource languages compared to English, corresponding to small (D=0.25D=-0.25) and medium (D=0.48D=-0.48) effect size.
We found a high degree of linguistic errors, particularly in non-English languages, where linguistic errors dominated the distribution of misrepresentations (see Figure 3). Moreover, the average number of logical errors was also higher for Indic languages. Overall, these findings suggested that models not only produce a high degree of cultural misrepresentations in Indic languages, but the overall quality of generated stories is poorer.
Resources Languages
High English
Mid Hindi, Bengali, Nepali, Tamil
Low Telugu, Marathi, Gujarati, Punjabi,
Kannada, Odia, Malayalam, Sindhi, Urdu
Table 6. Languages in which LLMs were evaluated, grouped by amount of resources available as per Chang et al. (2024).
Cultural inaccuracies were the most frequent category in English stories, averaging $0.8$ misrepresentations per story. Interestingly, this average reduces to $0.7$ and $0.4$ for mid and low resource languages. This can be attributed to lower number of CSIs in stories in mid and low resource languages. This difference was statistically significant ($p<0.001$), with the median number of CSIs per story decreasing by $8$ for mid-resource and $11$ for low-resource languages compared to English, corresponding to small effect sizes ($D=0.20$ and $D=0.19$, respectively).
Frequency of misrepresentation across tiers
Stratifying the misrepresentations based on the tier of the regions, we found that except for Aya 3232B, all other models generated more misrepresentations for tier-$3$ regions as compared to tier-$1$ regions. Using the Mann-Whitney test, we observed that this difference was statistically significant. Specifically, stories set in tier-$3$ regions contained one more misrepresentation per story compared to tier-$1$ ($p<.001$, D=0.21D=-0.21) and tier-$2$ regions ($p=.039$, D=0.11D=-0.11).
@@
We found that Gemini $2.5$ Pro had a higher number of CSIs, indicating cultural richness in the stories with low fractions of these CSIs being misrepresented. This indicated that, overall, it performs better at representing culture. Interestingly, we see that while LLaMA $3.3$ 7070B did not generate many misrepresentations per story, the generated stories exhibited the lowest cultural richness, as measured by CSI counts (Table 5). This highlighted the trade-off between generating stories with limited cultural representation and those containing cultural misrepresentations.
Next, we analyzed the CSI categories that are frequently misrepresented (see Figure 5 in the Appendix). We find that most cultural inaccuracies are around social practices, while factual errors tend to be about geographical facts. We also note a high degree of errors related to food items, but fewer errors related to topics like history and kinship. These findings highlight clear areas where cultural representations are lacking for improvement in future models.
@@
Relatability Ratings
Model-generated stories averaged a relatability rating of $3.25$ on a scale of $1$ to $5$ as presented in Table 5. This suggests a clear headroom for improvement. Of all models, Gemini $2.5$ Pro received the highest relatability rating, while open-source models generally lagged behind. Using a Mann–Whitney U test, we observed a statistically significant increase of one rating for closed-source models compared to open-source models ($p<.001$, D=0.24D=-0.24) with a small effect size.
5. Measuring Cultural Knowledge (RQ3)
Our third research question explores whether the observed cultural misrepresentations can be attributed to the models’ insufficient cultural knowledge of region-specific norms, values, artifacts, and expressions. To address this, we converted the span annotations of misrepresentations obtained in the human evaluations to standalone questions of cultural knowledge. We used GPT $4.1$ to generate candidate questions from the misrepresentation spans, and then human annotators verified each one of them. They refined the questions when necessary. This resulted in TALES-QA, a question bank consisting of 1,6001,600+ questions, which capture nuanced and diverse cultural knowledge about regions in India.
5.1. Question Bank Creation (TALES-QA)